OpenStudy (anonymous):

If circle P is tangent to both the x-axis and the y-axis and the coordinates of the center are (r,r) what is the slope of the line through the origin and the center?

OpenStudy (anonymous):

it's either y=x or y=-x.

OpenStudy (dumbcow):

m = (r-0)/(r-0) = r/r = 1

OpenStudy (anonymous):

y=x if in 1 or 3 quad, y=-x if in 2 or 4

OpenStudy (dumbcow):

oh yeah, prob has a picture that shows which quadrant usually its quad 1